# Hyperspectral Image Data Reduction for Endmember Extraction

### v1

@@ 1 Introduction
Endmember extraction from a hyperspectral image (HSI) is a fundamental task in utilizing hyperspectral imaging technology for Earth’s surface observation, with applications ranging from vegetation monitoring to mineral and resource exploration. Numerous endmember extraction methods have been proposed to date; see [Bio12, Ma14, Gil20] for surveys of recent advances. Among these, self-dictionary methods have received growing attention because of their ability to achieve high extraction accuracy on HSIs. However, a major drawback of self-dictionary methods is their high computational cost, which hinders their applicability to large-scale HSIs.
Self-dictionary methods formulate the endmember extraction task as a sparse optimization problem, treating the HSI data itself as a dictionary, and then solve a convex relaxation of this problem. Endmembers are identified using the optimal solution of the relaxed problem. Although various self-dictionary methods have been developed in the literature [Bit12, Elh12, Ess12, Gil18, Ngu22], this paper mainly focuses on the method proposed by Bittorf et al. [Bit12], which employs a linear programming (LP) formulation as the convex relaxation. For brevity, we refer to this LP-based self-dictionary method simply as the “LP method” throughout the paper. Solving this LP relaxation is the primary computational bottleneck in applying the LP method to large-scale HSIs, because the size of the LP problem grows quadratically with the number of pixels in the image.
There are two main approaches to addressing this computational issue: (i) developing efficient algorithms for solving the convex relaxation problem, and (ii) developing preprocessing techniques that reduce the size of the HSI data before applying self-dictionary methods. There is a line of research exploring the first approach. For example, an efficient algorithm based on the column generation framework was proposed in [Miz25] for the LP method. However, even with this improvement, processing the Urban HSI dataset, which is a widely used benchmark in the remote sensing community, still takes 8 hours. In light of this situation, this paper pursues the second approach.
@@
At this point, a natural question arises: After removing the columns that lie in the interior of the conical hull of the HSI matrix, how close are the remaining ones to the endmembers? To the best of our knowledge, this research question has not been investigated from a theoretical perspective in the existing literature. In the works of Esser et al. [Ess12] and Gillis and Luce [Gil18], data reduction techniques based on clustering were employed within self-dictionary methods. However, the theoretical properties of these reduction steps were not analyzed in their studies. In this paper, we address this question through a theoretical analysis. Building on the resulting insights, we then develop a data reduction algorithm, which we integrate into a self-dictionary method.
@@
Second, we develop a data reduction algorithm based on the theoretical result, and subsequently propose a data-reduced self-dictionary method, abbreviated as REDIC, for endmember extraction from HSIs by integrating data reduction with the variant of the LP method proposed in [Miz25]. As observed in the case of the Samson dataset, applying data reduction to HSI matrices can remove a large portion of their columns. However, directly applying the LP method to the reduced data does not necessarily yield high extraction accuracy. To address this issue, we incorporate a data augmentation process together with an averaging technique to improve the extraction accuracy of REDIC. For the reduced matrix obtained after data reduction, we augment it by randomly selecting some of the removed columns and adding them back. This augmentation step helps improve the accuracy of the extracted endmembers. However, the randomness involved in the augmentation process introduces variability into the extraction results. To mitigate this variability, we repeat the augmentation and endmember extraction using the LP method multiple times. Finally, we average the outputs produced by the LP method to obtain the final estimates of the endmembers.
Third, we present numerical experiments demonstrating the performance of REDIC. The experiments were conducted using both synthetic and real HSI datasets. The results indicate that the data reduction algorithm within REDIC can remove a large portion of the columns in HSI matrices; for example, for the Urban dataset, which originally contains about one hundred thousand columns, only 483 columns remain after applying this data reduction. In addition, among the remaining columns, there exists at least one column close to each of the endmembers. We compared the endmember extraction accuracy and computation time of REDIC with those of the LP method in [Miz25]. The results show that REDIC can achieve extraction accuracy comparable to that of the LP method while being substantially faster; for the Urban dataset, REDIC requires only 16 minutes, whereas the LP method takes 8 hours.
@@ 2.1 Problem Formulation
Problem 1 is essentially equivalent to computing nonnegative matrix factorizations (NMFs) under the separability assumption. Considerable research has been devoted to the separable NMF problem; see [Fu19, Gil20] for surveys of recent advances. Arora et al. [Aro12a] showed that $W$ can be recovered from $A$ in polynomial time when $V=O$. Even when $V\neq O$, it is still possible to extract a submatrix of $A$ that is close to $W$, provided that the noise $V$ is below a certain threshold. Separable NMFs have applications not only to endmember extraction from HSIs, but also to topic modeling in document corpora [Aro12b, Aro13]. Other applications include community detection [Hua19], spectral clustering [Miz22b], and crowdsourcing [Ibr19].
@@ 2.2 Review of Self-Dictionary Methods
Numerous endmember extraction methods have been proposed to date. These methods can be broadly classified into two categories: greedy methods and convex programming-based methods. Representative examples of greedy methods include PPI [Boa95], N-FINDR [Win99], SPA [Ara01, Gil14a], and VCA [Nas05], all of which are widely used for endmember extraction. Greedy methods have low computational cost and can therefore handle large-scale datasets, although the accuracy of the extracted endmembers may not always be sufficient. In contrast, convex programming-based methods generally achieve higher extraction accuracy, but their high computational cost remains a drawback.
@@
Solving problem (3) exactly is difficult because of the combinatorial constraint $\|X\|_{\mathrm{row},0}=r$. Self-dictionary methods therefore typically rely on a convex relaxation of this problem. As mentioned in Section 1, this paper focuses on the LP method proposed by Bittorf et al. [Bit12], which employs an LP formulation as the convex relaxation. This method was originally developed in the context of identifying hot topics in document corpora and was therefore named Hottopixx. Since its original proposal, the LP method has been extended and refined by several researchers [Gil13, Gil14b, Miz22a]. In what follows, we review the version of the LP method proposed in [Miz22a]. This method is based on the following optimization model:
@@
The LP method consists of two main steps: it first computes the optimal solution $X_{\mathrm{opt}}$ of $\mathsf{H}$ and then performs a postprocessing step based on $X_{\mathrm{opt}}$. In the postprocessing step, $r$ clusters of the column indices of the input matrix $A$ are formed according to the diagonal elements of $X_{\mathrm{opt}}$. One representative element is then chosen from each cluster and returned as the output. Theorem 3.2 of [Miz22a] guarantees that the corresponding columns of $A$ indexed by these elements are close to the $r$ endmember signatures, provided that the noise level is below a certain threshold.
From a theoretical perspective, the LP method is attractive, as the theoretical result suggests that the accuracy of the extracted endmembers can be high. However, its major drawback lies in the high computational cost. Specifically, the method requires solving the model $\mathsf{H}$, which reduces to an LP with $O(n^{2})$ variables and $O(n^{2})$ constraints; see [Miz25] for details. Consequently, the method becomes impractical for HSIs with a large number of pixels $n$. To address this computational issue, a row and column expansion (RCE) algorithm was proposed in [Miz25] to solve model $\mathsf{H}$ efficiently. The RCE algorithm is based on a column generation technique, a well-known approach for solving large-scale LPs. Nevertheless, even with RCE, the improvement in computation time is insufficient. For example, as shown in Section 5.3, the Urban dataset requires 8 hours of computation, whereas greedy methods such as SPA and VCA finish in only a few seconds. This observation motivates us to develop a data reduction technique for hyperspectral images.
In addition to the LP method, various self-dictionary approaches have been proposed for endmember extraction from HSIs [Elh12, Ess12, Gil18, Ngu22]. FGNSR, proposed by Gillis and Luce [Gil18], and MERIT, proposed by Nguyen et al. [Ngu22], are most closely related to the LP method. These methods are based on the following optimization model:
@@
where $\lambda$ is a penalty parameter and $\Phi$ is a regularization term designed to promote row sparsity in $X$. FGNSR and MERIT identify endmembers by leveraging the optimal solution of this model. As with the LP method, solving this model incurs a high computational cost and becomes impractical for HSIs with a large number of pixels. The work by Esser et al. [Ess12] can be viewed as a self-dictionary approach to the endmember extraction problem, whereas the work by Elhamifar et al. [Elh12] addresses a related problem within a similar modeling framework.
@@ 3.1 Main Theorem
A similar assumption has been used in the analyses of the noise robustness of LP methods for separable NMFs [Gil13, Miz22a]. While [Miz22a] assumes $0\leq\epsilon<1$, this condition is automatically satisfied in the setting of Theorem 1 due to the remaining assumptions. Part (a) can be imposed without loss of generality for $r$-separable matrices. Our analysis employs the conditioning parameter $\rho(W)$ for $W\in\mathbb{R}^{d\times r}_{+}$, defined by
@@ 3.2 Data Reduction Algorithms
We conclude this section with a discussion of related work. Finding a set $\mathcal{K}\in\Gamma(A)$ with the smallest number of elements is referred to as the conical hull problem in the paper by Dulá et al. [Dul98]. They proposed an efficient algorithm for solving this problem and reported numerical results. Starting with $\mathcal{K}=\emptyset$, their algorithm constructs $\mathcal{K}$ by sequentially adding elements. This stands in contrast to our DR algorithm, which begins with $\mathcal{K}=[n]$, where $n$ is the number of columns of $A$, and removes elements one by one. At each iteration, their algorithm performs a feasibility test using an LP formulation, whereas our DR algorithm employs a nonnegative least-squares formulation. As mentioned in [Fuk18], their algorithm can be regarded as an instance of the algorithmic framework studied in [Cla94].
@@ 4 The REDIC Method for Endmember Extraction
We develop REDIC, a data-reduced self-dictionary method for endmember extraction from HSIs by combining DRS with the LP method of [Miz22a]. Here, we describe the details of REDIC. To implement the LP method, REDIC employs the approach proposed in [Miz25]. This approach is referred to as EEHT (efficient and effective Hottopixx), where Hottopixx is the original name of the LP method.
@@
Here, $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix whose entries are the top $r$ singular values of $A$, and $U_{r}\in\mathbb{R}^{d\times r}$ (resp. $V_{r}\in\mathbb{R}^{n\times r}$) are the left (resp. right) singular vectors associated with these singular values. We then construct a reduced matrix $A^{\prime}=\Sigma_{r}V_{r}^{\top}\in\mathbb{R}^{r\times n}$. It should be noted that if $A$ follows the LMM, i.e., $A=WH+V$, then $A^{\prime}$ also follows the LMM, i.e., $A^{\prime}=W^{\prime}H+V^{\prime}$, but with $r$ rows instead of $d$ rows. Second, EEHT constructs the model $\mathsf{H}$ using $A^{\prime}$ instead of $A$ and solves it with the RCE algorithm proposed in [Miz25]. Finally, EEHT performs a postprocessing step: it partitions the column index set of $A$ into $r$ clusters based on the optimal solution of $\mathsf{H}$ and chooses elements from these clusters according to a prescribed criterion. For a detailed description of EEHT, we refer readers to Algorithm 4 in [Miz25]. The experimental results demonstrate that RCE significantly reduces the computation time required to solve $\mathsf{H}$. Moreover, the accuracy of the extracted endmembers can be improved when the postprocessing step chooses representative elements from clusters based on their centroids, a procedure referred to as method-C. EEHT employing method-C for postprocessing is denoted as EEHT-C in [Miz25].
@@ 5.1 Objectives and Experimental Setup
REDIC also performs RCE to solve the model $\mathsf{H}$. For the implementation of RCE, we used a reorganized and speed-improved version of the MATLAB code developed in [Miz25], in which CPLEX is employed as the LP solver. The comparison of computation time with the previous version of the RCE code is described in Remark 1. RCE first solves a small-scale subproblem of $\mathsf{H}$ and then gradually increases the problem size until the optimal solution of $\mathsf{H}$ is obtained. It includes the parameters $\zeta$ and $\eta$ to control the size of the initial subproblem. These parameter values were set as follows to maintain consistency with the experimental setup in [Miz25]: for an input matrix with $n$ columns, if $n\leq 300$, then $(\zeta,\eta)=(0,n)$; if $300<n\leq 50000$, then $(\zeta,\eta)=(10,100)$; and if $n>50000$, then $(\zeta,\eta)=(50,300)$.
We employed three real HSI datasets in our experiments: Jasper Ridge, Samson, and Urban, which are widely used for evaluating endmember extraction methods. These datasets were constructed following the procedure described in [Zhu17] and are the same as those used in [Miz25]. The details of each dataset are as follows:
@@
  • Urban: The Urban HSI was acquired by the HYDICE sensor and consists of $307\times 307$ pixels with 210 bands covering wavelengths from 400 nm to 2500 nm. We used the entire image with 162 valid bands after removing the degraded ones. This region contains four to six endmembers. Following the setting in [Gil18, Miz25], we set the number of endmembers to six: Asphalt, Grass, Tree, Roof 1, Roof 2, and Soil.

### v2

@@ 1 Introduction
Endmember extraction from a hyperspectral image (HSI) is a fundamental task in utilizing hyperspectral imaging technology for Earth’s surface observation, with applications ranging from vegetation monitoring to mineral and resource exploration. Numerous endmember extraction methods have been proposed to date; see [5, 22, 19] for surveys of recent advances. Among these, self-dictionary methods have received growing attention because of their ability to achieve high extraction accuracy on HSIs. However, a major drawback of self-dictionary methods is their high computational cost, which hinders their applicability to large-scale HSIs.
Self-dictionary methods formulate the endmember extraction task as a sparse optimization problem, treating the HSI data itself as a dictionary, and then solve a convex relaxation of this problem. Endmembers are identified using the optimal solution of the relaxed problem. Although various self-dictionary methods have been developed in the literature [6, 10, 11, 15, 27], this paper mainly focuses on the method proposed by Bittorf et al. [6], which employs a linear programming (LP) formulation as the convex relaxation. For brevity, we refer to this LP-based self-dictionary method simply as the “LP method” throughout the paper. Solving this LP relaxation is the primary computational bottleneck in applying the LP method to large-scale HSIs, because the size of the LP problem grows quadratically with the number of pixels in the image.
There are two main approaches to addressing this computational issue: (i) developing efficient algorithms for solving the convex relaxation problem, and (ii) developing preprocessing techniques that reduce the size of the HSI data before applying self-dictionary methods. There is a line of research exploring the first approach. For example, an efficient algorithm based on the column generation framework was proposed in [25] for the LP method. However, even with this improvement, processing the Urban HSI dataset, which is a widely used benchmark in the remote sensing community, still takes 8 hours. In light of this situation, this paper pursues the second approach.
@@
At this point, a natural question arises: After removing the columns that lie in the interior of the conical hull of the HSI matrix, how close are the remaining ones to the endmembers? To the best of our knowledge, this research question has not been investigated from a theoretical perspective in the existing literature. In the works of Esser et al. [11] and Gillis and Luce [15], data reduction techniques based on clustering were employed within self-dictionary methods. However, the theoretical properties of these reduction steps were not analyzed in their studies. In this paper, we address this question through a theoretical analysis. Building on the resulting insights, we then develop a data reduction algorithm, which we integrate into a self-dictionary method.
@@
Second, we develop a data reduction algorithm based on the theoretical result, and subsequently propose a data-reduced self-dictionary method, abbreviated as REDIC, for endmember extraction from HSIs by integrating data reduction with the variant of the LP method proposed in [25]. As observed in the case of the Samson dataset, applying data reduction to HSI matrices can remove a large portion of their columns. However, directly applying the LP method to the reduced data does not necessarily yield high extraction accuracy. To address this issue, we incorporate a data augmentation process together with an averaging technique to improve the extraction accuracy of REDIC. For the reduced matrix obtained after data reduction, we augment it by randomly selecting some of the removed columns and adding them back. This augmentation step helps improve the accuracy of the extracted endmembers. However, the randomness involved in the augmentation process introduces variability into the extraction results. To mitigate this variability, we repeat the augmentation and endmember extraction using the LP method multiple times. Finally, we average the outputs produced by the LP method to obtain the final estimates of the endmembers.
Third, we present numerical experiments demonstrating the performance of REDIC. The experiments were conducted using both synthetic and real HSI datasets. The results indicate that the data reduction algorithm within REDIC can remove a large portion of the columns in HSI matrices; for example, for the Urban dataset, which originally contains about one hundred thousand columns, only 483 columns remain after applying this data reduction. In addition, among the remaining columns, there exists at least one column close to each of the endmembers. We compared the endmember extraction accuracy and computation time of REDIC with those of the LP method in [25]. The results show that REDIC can achieve extraction accuracy comparable to that of the LP method while being substantially faster; for the Urban dataset, REDIC requires only 16 minutes, whereas the LP method takes 8 hours.
@@ 2.1 Problem Formulation
Problem 1 is essentially equivalent to computing nonnegative matrix factorizations (NMFs) under the separability assumption. Considerable research has been devoted to the separable NMF problem; see [12, 19] for surveys of recent advances. Arora et al. [2] showed that $W$ can be recovered from $A$ in polynomial time when $V=O$. Even when $V\neq O$, it is still possible to extract a submatrix of $A$ that is close to $W$, provided that the noise $V$ is below a certain threshold. Separable NMFs have applications not only to endmember extraction from HSIs, but also to topic modeling in document corpora [4, 3]. Other applications include community detection [20], spectral clustering [23], and crowdsourcing [21].
@@ 2.2 Review of Self-Dictionary Methods
Numerous endmember extraction methods have been proposed to date. These methods can be broadly classified into two categories: greedy methods and convex programming-based methods. Representative examples of greedy methods include PPI [7], N-FINDR [28], SPA [1, 16], and VCA [26], all of which are widely used for endmember extraction. Greedy methods have low computational cost and can therefore handle large-scale datasets, although the accuracy of the extracted endmembers may not always be sufficient. In contrast, convex programming-based methods generally achieve higher extraction accuracy, but their high computational cost remains a drawback.
@@
Solving problem (3) exactly is difficult because of the combinatorial constraint $\|X\|_{\mathrm{row},0}=r$. Self-dictionary methods therefore typically rely on a convex relaxation of this problem. As mentioned in Section 1, this paper focuses on the LP method proposed by Bittorf et al. [6], which employs an LP formulation as the convex relaxation. This method was originally developed in the context of identifying hot topics in document corpora and was therefore named Hottopixx. Since its original proposal, the LP method has been extended and refined by several researchers [17, 14, 24]. In what follows, we review the version of the LP method proposed in [24]. This method is based on the following optimization model:
@@
The LP method consists of two main steps: it first computes the optimal solution $X_{\mathrm{opt}}$ of $\mathsf{H}$ and then performs a postprocessing step based on $X_{\mathrm{opt}}$. In the postprocessing step, $r$ clusters of the column indices of the input matrix $A$ are formed according to the diagonal elements of $X_{\mathrm{opt}}$. One representative element is then chosen from each cluster and returned as the output. Theorem 3.2 of [24] guarantees that the corresponding columns of $A$ indexed by these elements are close to the $r$ endmember signatures, provided that the noise level is below a certain threshold.
From a theoretical perspective, the LP method is attractive, as the theoretical result suggests that the accuracy of the extracted endmembers can be high. However, its major drawback lies in the high computational cost. Specifically, the method requires solving the model $\mathsf{H}$, which reduces to an LP with $O(n^{2})$ variables and $O(n^{2})$ constraints; see [25] for details. Consequently, the method becomes impractical for HSIs with a large number of pixels $n$. To address this computational issue, a row and column expansion (RCE) algorithm was proposed in [25] to solve model $\mathsf{H}$ efficiently. The RCE algorithm is based on a column generation technique, a well-known approach for solving large-scale LPs. Nevertheless, even with RCE, the improvement in computation time is insufficient. For example, as shown in Section 5.3, the Urban dataset requires 8 hours of computation, whereas greedy methods such as SPA and VCA finish in only a few seconds. This observation motivates us to develop a data reduction technique for hyperspectral images.
In addition to the LP method, various self-dictionary approaches have been proposed for endmember extraction from HSIs [10, 11, 15, 27]. FGNSR, proposed by Gillis and Luce [15], and MERIT, proposed by Nguyen et al. [27], are most closely related to the LP method. These methods are based on the following optimization model:
@@
where $\lambda$ is a penalty parameter and $\Phi$ is a regularization term designed to promote row sparsity in $X$. FGNSR and MERIT identify endmembers by leveraging the optimal solution of this model. As with the LP method, solving this model incurs a high computational cost and becomes impractical for HSIs with a large number of pixels. The work by Esser et al. [11] can be viewed as a self-dictionary approach to the endmember extraction problem, whereas the work by Elhamifar et al. [10] addresses a related problem within a similar modeling framework.
@@ 3.1 Main Theorem
A similar assumption has been used in the analyses of the noise robustness of LP methods for separable NMFs [17, 24]. While [24] assumes $0\leq\epsilon<1$, this condition is automatically satisfied in the setting of Theorem 1 due to the remaining assumptions. Part (a) can be imposed without loss of generality for $r$-separable matrices. Our analysis employs the conditioning parameter $\rho(W)$ for $W\in\mathbb{R}^{d\times r}_{+}$, defined by
@@ 3.2 Data Reduction Algorithms
We conclude this section with a discussion of related work. Finding a set $\mathcal{K}\in\Gamma(A)$ with the smallest number of elements is referred to as the conical hull problem in the paper by Dulá et al. [9]. They proposed an efficient algorithm for solving this problem and reported numerical results. Starting with $\mathcal{K}=\emptyset$, their algorithm constructs $\mathcal{K}$ by sequentially adding elements. This stands in contrast to our DR algorithm, which begins with $\mathcal{K}=[n]$, where $n$ is the number of columns of $A$, and removes elements one by one. At each iteration, their algorithm performs a feasibility test using an LP formulation, whereas our DR algorithm employs a nonnegative least-squares formulation. As mentioned in [13], their algorithm can be regarded as an instance of the algorithmic framework studied in [8].
@@ 4 The REDIC Method for Endmember Extraction
We develop REDIC, a data-reduced self-dictionary method for endmember extraction from HSIs by combining DRS with the LP method of [24]. Here, we describe the details of REDIC. To implement the LP method, REDIC employs the approach proposed in [25]. This approach is referred to as EEHT (efficient and effective Hottopixx), where Hottopixx is the original name of the LP method.
@@
Here, $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix whose entries are the top $r$ singular values of $A$, and $U_{r}\in\mathbb{R}^{d\times r}$ (resp. $V_{r}\in\mathbb{R}^{n\times r}$) are the left (resp. right) singular vectors associated with these singular values. We then construct a reduced matrix $A^{\prime}=\Sigma_{r}V_{r}^{\top}\in\mathbb{R}^{r\times n}$. It should be noted that if $A$ follows the LMM, i.e., $A=WH+V$, then $A^{\prime}$ also follows the LMM, i.e., $A^{\prime}=W^{\prime}H+V^{\prime}$, but with $r$ rows instead of $d$ rows. Second, EEHT constructs the model $\mathsf{H}$ using $A^{\prime}$ instead of $A$ and solves it with the RCE algorithm proposed in [25]. Finally, EEHT performs a postprocessing step: it partitions the column index set of $A$ into $r$ clusters based on the optimal solution of $\mathsf{H}$ and chooses elements from these clusters according to a prescribed criterion. For a detailed description of EEHT, we refer readers to Algorithm 4 in [25]. The experimental results demonstrate that RCE significantly reduces the computation time required to solve $\mathsf{H}$. Moreover, the accuracy of the extracted endmembers can be improved when the postprocessing step chooses representative elements from clusters based on their centroids, a procedure referred to as method-C. EEHT employing method-C for postprocessing is denoted as EEHT-C in [25].
@@ 5.1 Objectives and Experimental Setup
REDIC also performs RCE to solve the model $\mathsf{H}$. For the implementation of RCE, we used a reorganized and speed-improved version of the MATLAB code developed in [25], in which CPLEX is employed as the LP solver. The comparison of computation time with the previous version of the RCE code is described in Remark 1. RCE first solves a small-scale subproblem of $\mathsf{H}$ and then gradually increases the problem size until the optimal solution of $\mathsf{H}$ is obtained. It includes the parameters $\zeta$ and $\eta$ to control the size of the initial subproblem. These parameter values were set as follows to maintain consistency with the experimental setup in [25]: for an input matrix with $n$ columns, if $n\leq 300$, then $(\zeta,\eta)=(0,n)$; if $300<n\leq 50000$, then $(\zeta,\eta)=(10,100)$; and if $n>50000$, then $(\zeta,\eta)=(50,300)$.
We employed three real HSI datasets in our experiments: Jasper Ridge, Samson, and Urban, which are widely used for evaluating endmember extraction methods. These datasets were constructed following the procedure described in [29] and are the same as those used in [25]. The details of each dataset are as follows:
@@
  • Urban: The Urban HSI was acquired by the HYDICE sensor and consists of $307\times 307$ pixels with 210 bands covering wavelengths from 400 nm to 2500 nm. We used the entire image with 162 valid bands after removing the degraded ones. This region contains four to six endmembers. Following the setting in [15, 25], we set the number of endmembers to six: Asphalt, Grass, Tree, Roof 1, Roof 2, and Soil.